A valószínűségszámítás és matematikai statisztika gyakorlaton való részvétel

dátum:  IX. 5. IX. 12. IX. 19. IX. 26. X. 3. X.10.

diákok:

Andrikó Imre  +  +  −  −  −  +
Brozsó András  +  +  +  +  −  +
Brumár László  +  +  +  −  +  +
Földi Péter  +  +  +  +  +  +
Guta Gábor  +  −  −  +  −  −
Gyulai Ágoston  +  +  +  +  +  +
Horváth Tibor  +  +  +  +  +  +
Krakkó Zoltán  +  +  +  +  +  +
Nádasi Péter  +  +  +  +  +  +
Péter Szabó Ramóna  +  +  +  +  +  +
Piros Attila  +  +  +  +  +  +
Podlovics Szilvia  +  +  +  +  +  +
Simon Imre  +  +  +  +  +  +
Szabó Lajos Tibor  +  +  −  +  +  +
Szabó Máté  +  +  +  +  +  +
Száldóbágyi Lóránd Lajos +  +  +  +  +  +
Szivák Gyula  +  +  +  +  +  +
Szlezák Gergely  +  +  +  +  +  +
Szlávik Linda  +  +  +  +  +  +
Szőke Péter Pál  +  +  +  +  +  +
Tordai Rita  +  +  +  +  −  +
Tóth Tamás  +  +  +  +  +  +
Török Norbert  +  +  +  +  +  −
Turják Balázs  +  +  +  +  +  +
Benedek István  +
Gyüre Zsolt  +
Nyircsák Péter  +

dátum:  X. 31. XI. 7. XI. 14. XI. 21. XI. 28. XII. 5. XII. 12. XII. 19.

diákok:

Andrikó Imre  +  +  +  +
Brozsó András  +  +  +  +
Brumár László  +  −  −  −
Földi Péter  +  +  +  +
Guta Gábor  −  −  −  +
Gyulai Ágoston  +  +  +  −
Horváth Tibor  +  +  +  +
Krakkó Zoltán  −  +  −  +
Nádasi Péter  +  +  +  +
Péter Szabó Ramóna  +  −  +  +
Piros Attila  +  +  +  −
Podlovics Szilvia  +  +  +  +
Simon Imre  +  +  +  +
Szabó Lajos Tibor  +  −  −  +
Szabó Máté  +  +  +  +
Száldóbágyi Lóránd Lajos +  +  +  +
Szivák Gyula  +  +  +  +
Szlezák Gergely  +  +  +  +
Szlávik Linda  +  +  +  +
Szőke Péter Pál  +  +  +  +
Tordai Rita  +  +  +  +
Tóth Tamás  +  +  +  +
Török Norbert  +  −  +  +
Turják Balázs  +  +  −  +